OpenStudy (anonymous):

In related rates how would you do this g(x)=(y(x))^3 , (1,3)

OpenStudy (amistre64):

the outer function is u^3 and it depends on that value of y(x): dg/dy y(x) is the next function, and it depends on the value of x: dy/dx x is the innermost function and drives the whole thing, it is either independant or relies on an external source such as time... dx/dx or dx/dt since x = 1; y(1) = 3 since y(1) = 3; g(x) = 3^3 = 27

OpenStudy (amistre64):

dg/dt = dg/dy* dy/dx* dx/dx (or dx/dt)